Water flows with speed v1 inside a long pipe and exits from the right into the air with a speed v2= 6.8m/s. The cross-sectional area of the opening on the right is 6 cm^2, while the rest of the pipe has a cross-section of 12 cm^2. The vertical standing column of water has height of 4.8 m. What is the elevation in meters from the base of the pipe to the opening. Pwater= 1000 kg/m^3
Answer #1

6.8 m/s Val A = 6 ay Them 4 = - lzen h=4.8 from equation of Continuty y (4) = b (4) V₂ (A2 Y(4) - 16:24 V = 3.4 m/s² from Bey(3,2) = yeh-he (4-32-6432 = 219086, the 4,- 1.76 h = 1.76 the 176 + 418 lugs. 9 :